# Efficient Nearest Neighbor Search Using Dynamic Programming

### v4

@@ IV-C Farthest Point Sampling
Figure 3: Coupling of Farthest Point sorting with incremental Voronoi Construction. Left: Uninserted points (gray) are distributed within the Voronoi cells of the inserted points (black). Right: When a new point is inserted, only the distance of the uninserted points within the green and yellow cells may need to be updated.
IV-D Incremental Delaunay Construction
Currently, numerous incremental methods for Delaunay construction, such as parallel acceleration [24, 25] and pre-sorting points [26], are commonly used to speed up the process. However, due to the specific requirements of farthest point sorting and the need to generate a query table during the construction phase in our approach, traditional acceleration techniques for Delaunay construction are not directly applicable.
During the incremental Delaunay construction process, one of the most time-consuming operations is locating the tetrahedron that contains a newly inserted point. According to Section IV-A and IV-B of our algorithm, when inserting point $p_{k}$, we have already constructed a query table for finding the nearest point within the set $\{p_{i}\}_{i=1}^{k-1}$. This query table stores information that allows for rapid identification of nearest points among the already inserted points. Leveraging this structure, we can quickly determine the nearest point to $p_{k}$. This nearest point then serves as the starting point for efficiently locating the tetrahedron containing $p_{k}$, significantly accelerating the Delaunay construction process.
IV-E Complexity Analysis
In this section, we will analyze the complexity of the farthest point sampling as a standalone component, rather than as part of the construction phase.
Time Complexity of Query Table Construction. The complexity of the Query Table Construction is the same as that of the incremental Delaunay construction. On average, the time complexity for constructing a Delaunay in 3D is $O(n\log n)$, while the worst-case time complexity is $O(n^{2})$. However, this worst-case scenario is seldom encountered in typical construction processes.
Space Complexity of the Query Table. On average, the number of neighbors for a Voronoi cell in 3D is a constant, denoted as $c$. This implies that when a new point is inserted, it is adjacent to $c$ cells on average, requiring the point to be appended to the end of $c$ Query Lists. Therefore, the average space complexity of the Query Table is
$O(\sum_{i=2}^{n}c)=c\times O(n)=O(n),$ (6)
where $n$ is the number of points.
Time Complexity of Nearest Neighbor Query. When querying the nearest neighbor for the point $q$, the complexity of the query phase depends on the number of times the cell containing $q$ (which may change during the query process) shrinks during the incremental construction, as a comparison is needed each time the Voronoi cell containing $q$ shrinks. Simply put, assuming $\Phi_{m}(q)$ is known and a new point $p_{m+1}$ is inserted, the cell of $p_{m+1}$ typically adjacent to $c$ cells on average. Therefore, the probability that $\text{Cell}(\Phi_{m}(q);\mathcal{V}_{m})$ shrinks is approximate $c/m$, indicating the possibility of needing a comparison. Thus, the average time complexity of the nearest neighbor query phase is
$O(\sum_{i=1}^{n-1}\frac{c}{i})=c\times O(\log n)=O(\log n).$ (7)
Time Complexity of Farthest Point Sampling. When we have already inserted $k$ points to construct the Voronoi diagram, each cell, on average, contains ${(n-k)}/{k}$ remaining points. Given the distances from all remaining points to the nearest sampled points are known, selecting the next point for sampling has a complexity of $O(\log(n-k))$. Knowing that the newly inserted point is adjacent to $c$ cells, the number of remaining points that potentially need their distances updated is $c\times(n-k)/k$.
Thus, the additional complexity caused by the farthest point sampling process is:
$O(\sum_{k=1}^{n-1}\left[\log(n-k)+\left(c\times\frac{n-k}{k}\right)\right])$
$=O(\sum_{k=1}^{n-1}\log(k)+cn\times\sum_{k=1}^{n-1}\frac{1}{k}-(n-1)\times c)$
$=O(n\log n)$ (8)
@@ V Results
Hardware environment. Our algorithm was implemented in C++ on a platform featuring a 3.4 GHz AMD Ryzen 9 5950X 16-Core CPU, 64GB of RAM, and running the Windows 11 operating system.
Experimental parameters. We utilized CGAL [19] for Delaunay computations. In our experiments, we used random sorting and farthest point sorting as insertion order. We measured and recorded both the construction performance and query performance for each method. The space for randomly generating query points was confined within the bounding box with 8x the volume (with edge lengths twice that of the original bounding box) of the input point cloud. The number of query points generated was set at 1000K.
Relevant approaches for comparison. Currently, there are numerous algorithms [27, 28, 29, 12] available for nearest neighbor search. In this paper, we selected two of the fastest and most classic methods for comparison.
@@
    KD Tree: We utilized the implementation from the nanoflann library [30], setting the maximum number of leaf nodes to 10. According to tests referenced in [31], this implementation is currently the fastest for nearest neighbor searches.
@@
    R Tree: We used the implementation from Boost Library [32].
In all the experiments below, Ours1 refers to our method using a random order for incremental Delaunay construction, while Ours2 indicates that the farthest point sorting strategy was applied.
V-A Query Table
Length of the query list for a vertex
Recall that the Query Table maintains a list of points encroached upon during the incremental construction of the Delaunay for each vertex in $V$. It is evident that the Voronoi cells of points inserted earlier often shrink more times, resulting in longer Query Lists. Therefore, to investigate how the length of the Query List varies with different insertion indexes, we conducted a study using the $100$K-vertex dragon point cloud as an example. As shown in Figure 4, we have plotted the length of the query list for each point with respect to its insertion order. Although the impact of random sorting and farthest point sorting strategies on the length of the Query List is not readily apparent from the figure, we calculated the average lengths of the Query Lists to be 15.684 and 14.165, with standard deviations of 17.705 and 16.491, respectively. The farthest point sorting resulted in a shorter and more uniform distribution of the Query List. Similar results were obtained across multiple data. This characteristic enables the query table generated through farthest point sorting to facilitate faster nearest-neighbor searches.
@@ V-A Query Table
Figure 4: The length of the Query List for different index points. Top: random sorting. Bottom: farthest point sorting.
Number of tested vertices
As shown in Algorithm 2, the nearest neighbor search problem is transformed into the task of calculating distances between points.
The number of distance comparisons between points during the query process partially reflects the efficiency of the nearest neighbor search. We tested several classical models and recorded the number of comparisons, along with the mean and variance of the Query Lists, as shown in Table I. It can be observed that, after utilizing the farthest point sorting method, both the average length (Avg) and variance (Var) of the Query Lists decreased, accompanied by a reduction in the average number of comparisons during the nearest neighbor search process.
@@
$\text{Ours}^{1}$ Avg 15.913 15.805 15.684 15.397 15.832 16.034 14.631
Var 17.911 17.264 17.705 18.834 17.631 17.968 20.431
Tested 171.783 166.726 209.516 182.719 234.742 276.549 418.285
$\text{Ours}^{2}$ Avg 14.414 14.345 14.165 13.964 14.370 14.478 13.067
Var 16.783 15.906 16.491 17.126 16.172 16.584 19.681
Tested 144.754 135.831 161.822 155.806 197.042 235.085 358.413
V-B Preprocessing Cost
The preprocessing time of our algorithm is primarily concentrated in the Delaunay construction phase, as it involves significant computational overhead in building and managing the complex graph structure. In this section, we present the preprocessing times of different algorithms across several classic models in Table II. Additionally, Figure 5 illustrates the preprocessing time relative to the resolution of the Dragon point cloud, providing a comparative analysis between our method and others. Furthermore, Figures 9 and Figure 10 show the preprocessing time of different methods on subsets of the ABC dataset [33] and the Thingi10k dataset [34]. It can be observed that at this stage, our method currently exhibits longer preprocessing times compared to existing nearest neighbor search algorithms, particularly at higher resolutions, indicating areas that require further optimization.
TABLE II: Preprocessing time ($s$) on classical models.
@@ V-B Preprocessing Cost
KD-tree 0.004 0.011 0.016 0.051 0.149 0.215 0.281
R-tree 0.005 0.014 0.022 0.062 0.170 0.243 0.311
Ours1 0.188 0.703 0.984 3.531 10.359 15.219 18.985
Ours2 0.265 0.765 1.406 4.782 15.157 22.953 26.938
@@
Figure 5: The preprocessing cost of KD Tree, R Tree and Ours on the Dragon data with varying resolutions. Top: preprocessing time. Bottom: the comparison about the preprocessing cost.
@@
Figure 6: Comparison about preprocessing cost on ABC dataset. Top: the time cost of preprocessing for KD Tree, R Tree and ours. Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.
@@
Figure 7: Comparison about preprocessing cost on Thingi10K dataset. Top: the time cost of preprocessing for KD Tree, R Tree and ours. Bottom: the comparison about the preprocessing cost among KD Tree, R Tree and ours.
V-C Query Performance
Using the Dragon dataset as an example, we present the query cost of our method compared to other approaches with respect to resolution in Figure 8, demonstrating the significant advantage of our method in this scenario. In Table III, we provide more detailed examples for several classical models.
[…]
Figure 8: Average query time on the Dragon model with varying resolutions. Top: Average query time ($\mu$s) among KD tree, R tree and ours. Bottom: The comparison about the average time different methods. Dashed line indicates where the ratio equals 1.
TABLE III: Average query time ($\mu s$) of KD Tree, R Tree, and Ours on classical models. The highest and second-fastest query times are highlighted in bold and underlined, respectively.
@@ V-C Query Performance
KD Tree 1.639 3.795 2.475 8.475 12.81 9.796 4.407
R Tree 1.623 2.739 2.034 5.359 5.241 4.518 1.851
Ours1 0.423 0.559 0.540 0.957 1.132 1.293 2.146
Ours2 0.344 0.482 0.388 0.752 0.834 1.070 1.694
V-D Tests on ABC & Thingi10K
Thingi10K [34] and ABC [33] are two large-scale 3D datasets that contain diverse models. To conduct a comprehensive comparison among KD Tree, R Tree, and Ours, we ran them on the dataset to compare both the preprocessing cost and the query cost. The query time are shown in Figures 9 and Figure 10. And the preprocessing time can be found in Figure 6 and Figure 7.
@@ V-D Tests on ABC & Thingi10K
Figure 9: Comparison about query performance on ABC dataset. Top: the average timing cost per query ($\mu$s) for KD Tree, R Tree and Ours. The y-axis is scaled logarithmically. Bottom: the comparison about the query cost among KD Tree, R Tree and Ours. Dashed line indicates where the ratio equals 1.
@@
Figure 10: Comparison about query performance on Thingi10K data. Top: the average timing cost per query ($\mu$s) for KD Tree, R Tree and Ours. The y-axis is scaled logarithmically. Bottom: the comparison about the query cost among KD Tree, R Tree and Ours. Dashed line indicates where the ratio equals 1.
V-E Other Types of Data
Uniform Point Cloud.
To fully demonstrate the efficiency of the algorithm, we compared KD Tree, R Tree, and Ours on uniform point clouds. As shown in Figure 11, it can be observed that even on completely uniform point clouds, our method remains comparable in query efficiency to the fastest current implementation of KD Tree and R Tree.
@@ V-E Other Types of Data
Figure 11: Average query time on the uniform point cloud with varying resolutions. Top: Average query time per query ($\mu$s) among KD Tree, R Tree and Ours. Bottom: The comparison of the average time of different methods. Dashed line indicates where the ratio equals 1.
Earth City.
Additionally, we tested the efficiency of the KD Tree, R Tree, and our method on an Earth data, distributed over a sphere, with the results shown in Figure 12. At this distribution, both KD Tree and R Tree exhibit a significant drop in efficiency. However, our method remains extremely fast, achieving up to 40x the query speed of KD Tree.
@@
Figure 12: Average query time on the earth data with varying resolutions. Top: Average query time among KD tree, R tree and ours. Bottom: The comparison about the average time different methods. Dashed line indicates where the ratio equals 1.
V-F Near and far query points.
Additionally, we observed that the location of query points at different positions can impact the efficiency of the query. To further investigate this, we conducted tests on uniformly distributed point clouds, point clouds representing Earth cities, and the Dragon model, with the number of points in each point cloud controlled at 1, [^0] K. The results are presented in Figure 13. The horizontal axis represents the ratio of the length of the generation space for the query points to the length of the bounding box of the target points along any coordinate axis. It can be observed that even in a completely uniform point cloud distribution, where the query point generation space matches the bounding box of the target points, our query speed remains comparable to that of the KD Tree. In other scenarios, our method demonstrates a significant advantage in query speed.

### v5

@@ IV-E Time Complexity Analysis
Figure 5: Representative 3D models from the ABC dataset [25].
[…]
Figure 6: Representative 3D models from the Thingi10K dataset [26].
[…]
Figure 7: Representative 3D models from the MedShapeNet dataset [27].
@@ V Results
Platform. Our algorithm was implemented in C++ and evaluated on a Mac mini equipped with an M4 CPU, 16GB of RAM, running the macOS operating system. We utilized CGAL [18] for Delaunay triangulation computations and employed CGAL’s multi-scale sorting approach and farthest point sorting as the insertion order strategies.
Relevant approaches for comparison. We selected six classic methods as comparison baselines, and together with our proposed method (incorporating two different pre-sorting strategies), we evaluated a total of eight approaches.
@@
    KD Tree: From the nanoflann library [28], with the maximum number of leaf nodes set to the default value of 10, which achieves nearly optimal query performance according to the benchmark results provided by the official repository.
@@
    R* Tree: From Boost library [29], with the maximum number of elements per node set to 10, consistent with the KD Tree configuration.
  3. 3.
    Octree: From PCL library [30], with the resolution parameter set to 3 times the inter-point distance.
  4. 4.
    BD Tree: From ANN library, with the maximum number of elements in leaf node set to 10, and we only use it for exact nearest neighbor queries.
  5. 5.
    Voronoi: From CGAL library, utilizing its provided nearest_vertex function for nearest neighbor queries.
  6. 6.
    Grid: Our own implementation, with cell side length set to 3 times the inter-point distance.
  7. 7.
    Ours1: Our method, using CGAL’s multi-scale sorting method spatial_sort for incremental Delaunay construction.
  8. 8.
    Ours2: Our method, using farthest point sorting strategy for incremental Delaunay construction.
Additionally, it is worth noting that the Grid method exhibits slower performance when there is a considerable distance between query points and target points. The Voronoi-based method is also relatively slow. Therefore, we only include results from these two methods in certain experiments.
Datasets and Query Point Generation. We selected three datasets and several standard models for algorithm evaluation. For nearest neighbor queries, we utilize the vertices of these 3D models as the point cloud data. After filtering out models that are excessively large or small, the datasets cover the following ranges:
  1. 1.
    ABC Dataset [25]: We tested on a subset of 10,000 models, with the final dataset covering 600 to 600,000 points. Figure 5 shows representative models from this dataset.
  2. 2.
    Thingi10K Dataset [26]: Originally containing 10,000 models, the filtered dataset covers 400 to 140,000 points. Figure 6 displays sample models from this collection.
  3. 3.
    MedShapeNet Dataset [27]: We evaluated on a subset of 10,000 models, with the processed dataset covering 200 to 150,000 points. Figure 7 illustrates representative models from this dataset.
For query point generation, we randomly distribute 1 million query points within the 2× axis-aligned bounding box for each test case and perform nearest neighbor queries.
@@
Figure 8: The length of the Query List for different index points. Top: CGAL’s multi-scale sorting. Bottom: farthest point sorting.
[…]
Figure 9: Query process convergence under two insertion strategies. Average nearest distance (normalized by point cloud diagonal) versus number of comparisons, averaged over one million queries.
@@
$\text{Ours}^{1}$ Avg 16.753 16.445 16.371 15.508 15.864 16.106 15.765
Var 20.099 19.281 16.619 19.181 19.026 19.301 18.713
Tested 188.0 183.96 240.15 214.39 274.27 304.22 339.05
$\text{Ours}^{2}$ Avg 14.414 14.340 14.197 14.115 14.246 14.333 13.638
Var 16.783 15.794 16.355 16.207 15.947 16.016 15.927
Tested 144.87 136.28 162.68 152.35 193.15 211.39 235.85
[…]
Figure 10: Preprocessing time comparison across ABC, Thingi10K, and MedShapeNet datasets (from left to right).
[…]
Figure 11: Preprocessing cost comparison of various methods on the Dragon model across different resolutions. Top: absolute preprocessing time. Middle and Bottom: relative preprocessing cost ratios between methods.
V-A Query Table
Analysis of Query List Length. Recall that each vertex maintains a Query List, which is a vector storing the subsequently inserted points that become adjacent to the current vertex in the Delaunay triangulation during the incremental construction process. It is evident that points inserted earlier experience more subsequent insertions, resulting in longer Query Lists as they accumulate more neighbors over time. Therefore, to investigate how the length of the Query List varies with different insertion indexes, we conducted a study using the 100100100K-vertex dragon point cloud as an example. As shown in Figure 8, we have plotted the length of the query list for each point with respect to its insertion order. The results reveal that two different insertion orders lead to distinctly different query length distribution trends. CGAL’s multi-scale sorting method exhibits a clear two-stage characteristic: the first stage demonstrates significantly longer average query lengths compared to the second stage, while the query lengths fluctuate around their respective average values within each stage. In contrast, the query lengths obtained through farthest point sampling show a steady declining trend.
Number of Comparisons. As shown in Algorithm 2, the nearest neighbor search problem is transformed into the task of calculating distances between points. The number of distance comparisons between points during the query process partially reflects the efficiency of the nearest neighbor search. We tested several classical models and recorded the number of comparisons, along with the mean and variance of the Query Lists, as shown in Table I. It can be observed that, after utilizing the farthest point sorting method, both the average length (Avg) and variance (Var) of the Query Lists decreased, accompanied by a reduction in the average number of comparisons during the nearest neighbor search process.
Query Process Convergence. In accordance with Algorithm 2, our method iteratively evaluates target points during query execution while dynamically updating the current nearest distance estimate. To quantitatively assess the influence of varying insertion orders on query efficiency, we conducted a comprehensive evaluation using the 100K dragon model. Specifically, we generated 1 million query points within the search space and executed nearest neighbor queries for each point. During execution, we systematically recorded the current nearest distance at each comparison step, subsequently computing the average distance across all queries per step. Figure 9 demonstrates the convergence behavior of this average nearest distance (normalized as a percentage of the point cloud’s diagonal length) as a function of comparison count under two distinct insertion ordering strategies. The results reveal that the Farthest Point Sampling strategy facilitates substantially accelerated and more consistent convergence throughout the query process, empirically validating its efficacy in enhancing search performance.
TABLE II: Preprocessing time ($s$) on classical models
@@ V-A Query Table
KD-tree 0.003 0.006 0.009 0.025 0.068 0.098 0.123
R-tree 0.004 0.007 0.010 0.028 0.078 0.109 0.170
Octree 0.002 0.003 0.004 0.011 0.033 0.047 0.055
BD-tree 0.005 0.010 0.014 0.044 0.118 0.163 0.198
Grid 0.002 0.004 0.008 0.024 0.313 0.362 0.053
Voronoi 0.057 0.143 0.189 0.517 1.359 1.943 2.466
Ours1 0.072 0.182 0.249 0.685 1.736 2.481 3.170
Ours2 0.103 0.327 0.466 1.698 5.005 7.454 9.602
V-B Preprocessing Cost
The preprocessing time of our algorithm is primarily concentrated in the Delaunay construction phase, as it involves significant computational overhead in building and managing the complex graph structure. In this section, we present the preprocessing times of different algorithms across several classic models in Table II. Figure 11 further illustrates how preprocessing time scales with point cloud resolution using the Dragon model as a representative example. We extend this evaluation to larger datasets in Figure 10, which demonstrates preprocessing performance across the ABC [25], Thingi10k [26], and MedShapeNet [27] datasets. It can be observed that at this stage, our method currently exhibits longer preprocessing times compared to existing nearest neighbor search algorithms, particularly at higher resolutions, indicating areas that require further optimization.
V-C Query Performance
Query efficiency constitutes the primary advantage of our proposed approach. To systematically evaluate this performance benefit, we conducted a multi-faceted analysis across various datasets and resolutions.
Using the Dragon dataset as a controlled benchmark, we measured query costs with respect to point cloud resolution, as illustrated in Figure 12. This analysis demonstrates a consistent and significant performance advantage for our method across all tested resolutions. Table III complements this analysis by providing detailed query performance comparisons across a diverse set of classical models at their standard resolutions, further validating our method’s superior efficiency.
To establish the generalizability of these performance advantages, we expanded our evaluation to large-scale datasets with diverse geometric characteristics: Thingi10K [26], ABC [25], and MedShapeNet [27]. Figure 13 summarizes these comprehensive results, revealing consistent performance patterns across all tested repositories. Notably, our method achieves query acceleration factors of 1-10× compared to current state-of-the-art approaches, with substantial improvements observed across various model types.
V-D Memory Usage
Space complexity represents another critical dimension for algorithm evaluation. To provide a comprehensive assessment, we analyzed the memory requirements of our approach alongside existing methods. The primary memory consumption in our algorithm stems from three components: incremental Delaunay triangulation construction, Query Table storage, and segment tree structures when implementing farthest point sampling.
It should be noted that precise memory quantification presents inherent challenges due to system-level memory management mechanisms, resulting in measurement fluctuations. For statistical robustness, our analysis excludes data points with memory utilization below 0.2M. Figure 14 presents the memory consumption of our algorithm across the ABC, Thingi10K, and MedShapeNet datasets.
The empirical results demonstrate that while our method exhibits moderately increased memory requirements compared to traditional approaches, this additional memory overhead remains within acceptable bounds given the significant query performance advantages. This trade-off between space complexity and query efficiency represents a favorable compromise for applications where query speed is paramount.
@@ V-D Memory Usage
Figure 12: Average query time on the Dragon model across varying resolutions. Top: Absolute query time ($\mu$s) for different methods. Middle and Bottom: Performance ratios between methods, with dashed line indicating equal performance (ratio = 1).
@@
Figure 13: Query time comparison across ABC, Thingi10K, and MedShapeNet datasets (from left to right). Dashed line indicates equal performance (ratio = 1).
@@
Figure 14: Memory utilization comparison across ABC, Thingi10K, and MedShapeNet datasets.
TABLE III: Average query time ($\mu$s) comparison across classical models. The highest and second-fastest query times are highlighted in bold and underlined, respectively
@@
KD-tree 1.109 2.542 1.585 5.402 6.003 4.296 1.815
R*-tree 1.450 2.584 1.808 4.974 4.967 3.616 1.446
Octree 5.060 8.475 5.719 17.06 17.21 14.89 10.22
BD-tree 2.344 4.721 5.715 13.11 78.04 23.02 2.674
@@
Voronoi 6.815 9.965 5.692 27.01 26.48 39.10 36.92
@@
Ours2 0.201 0.235 0.213 0.322 0.377 0.401 0.617
V-E Performance on Uniform Distributions
To evaluate our algorithm’s efficiency across different data distributions, we tested performance on uniform point clouds. Unlike previous experiments, we generated query points within exactly the same spatial bounds as the target points—not in an expanded bounding box. As shown in Figure 15, which presents both query and preprocessing times, our method achieves query efficiency comparable to optimized KD Tree and Grid implementations while outperforming other approaches. These results demonstrate our algorithm’s effectiveness across diverse point distribution patterns beyond its primary optimization for manifold data.
@@ V-E Performance on Uniform Distributions
Figure 15: Performance evaluation on uniform point clouds with varying numbers of target points. Top: Query time comparison shows our method’s efficiency comparable to KD Tree and Grid while outperforming others. Bottom: Preprocessing time comparison. All query points were generated within the exact same spatial bounds as the target points.
V-F Near and Far Query Points.
The spatial relationship between query points and target points significantly influences nearest neighbor search efficiency. To quantify this effect, we conducted experiments varying the query space extent relative to the target point cloud. We tested three distinct distributions: uniform point clouds, Earth city models, and the Dragon dataset, each containing 1 million points.
Figure 16 presents query performance as a function of spatial scaling. The horizontal axis represents the ratio between query generation space and target point bounding box (e.g., a value of 2 indicates query points generated within a 2× axis-aligned bounding box). Corresponding preprocessing times are provided in Table IV.
The results reveal an interesting pattern: even in uniform distributions, where our method shows comparable performance when query points share the same bounds as target points, our approach gains significant advantages as soon as the query space extends slightly beyond the target bounds. For non-uniform distributions like Earth city models and Dragon dataset, our method consistently outperforms traditional approaches across all spatial configurations.
@@
Figure 16: Average query time with varying sampling box sizes. The horizontal axis shows the scaling factor applied to each dimension of the target point bounding box to define the query generation space. Results for uniform point cloud (top), Dragon model (middle), and Earth city data (bottom).
TABLE IV: Preprocessing time comparison for uniform point cloud, Dragon model, and Earth city data (1M points each).
Octree BD Tree R* Tree KD Tree Ours1 Ours2
Uniform 0.047 0.145 0.121 0.105 2.317 6.775
Dragon 0.053 0.169 0.109 0.094 2.483 7.092
Earth 0.051 0.185 0.102 0.104 2.232 7.139
VI Application
We validate the effectiveness of our algorithm through five applications.
VI-A Point-to-Mesh Distance Queries
Given a mesh surface, fast querying of the closest geometric primitive (vertex, edge, or face) to a user-specified point, along with the corresponding closest point and minimum distance, represents a fundamental operation across diverse research domains [31, 32, 33, 34] including computer graphics, physical simulation, computational geometry, and computer-aided design. P2M [3] currently stands as the fastest point-to-mesh distance query algorithm. During preprocessing, P2M constructs Voronoi diagrams of model point clouds and builds interception tables for each point cloud’s Voronoi cell. In the query phase, it first employs a KD-tree to compute the nearest model vertex, then utilizes the interception table to obtain the nearest distance to the mesh.
In our experiments, we replace the KD-tree component with our algorithm. Since P2M already requires Voronoi diagram construction, our data structure introduces minimal additional preprocessing overhead while achieving 1-5× query speedup. We validate these speed improvements on the ABC dataset and several standard models, comparing against BVH-based implementations from Libigl [35] and FCPW [36] libraries (with vectorization enabled), as well as the original P2M method. Following the experimental setup of the P2M paper, for each test model, we randomly sample a million query points within the 10× axis-aligned bounding box. Query time results are presented in Figure 17 and Table V. The modified P2M algorithm demonstrates significant efficiency gains, achieving 1-5× speedup over the original P2M query performance. Preprocessing times are shown in Figure S1 and Table S1 of the supplementary material, where our preprocessing overhead remains nearly identical to the original P2M implementation.